OpenStudy (anonymous):

Given below is the odd-number function. Which of the following are equal to O(24)? Check all that apply O(n) = 2n - 1 a. the 24th odd number b. O(23)+1 c. 24 d. 47

OpenStudy (unklerhaukus):

d. and a.

OpenStudy (anonymous):

thank you!

OpenStudy (unklerhaukus):

can you see why?

OpenStudy (anonymous):

no, i was having a hard time with this problem, thats why i posted it

OpenStudy (unklerhaukus):

The odd numbers are all results of this function\[O(n)=2n-1\] The first odd number is \[O(1)=2(1)-1=1\] the second odd number is \[O(2)=2(2)-1=3\] so the 24th odd number is \[O(24)=2(24)-1=48-1=47\]. The next odd number after the nth odd is\[O(n+1)=2(n+1)-1=2n+2-1=-1=2n-1+2=O(n)+2\]

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

given below is the odd-number function. which of the following are equal to o(5)

OpenStudy (unklerhaukus):

below.. .